OpenStudy (anonymous):

a)The difference between 2 numbers is less than 1. Can the differnce between their squares be greater than 100? b)Each side of a rectangle became 1ft longer. Can its area become 100 sqfeet greater? Problems a) and b) are related. How?

ganeshie8 (ganeshie8):

@KingGeorge

ganeshie8 (ganeshie8):

@myininaya

OpenStudy (anonymous):

a) Call a and b the 2 numbers. we have: a - b < 1 a^2 - b^2 > 100 a^2 - b^2 = (a - b)(a + b) > 100 --> (a + b) >100/(a - b). If the sum of the 2 numbers is > 100, then their squares will be > 100. b) Call a and b the 2 sides of the rectangle. The area of the rectangle gets 100 sq ft bigger: (a + 1)(b + 1) > ab + 100 ((a + 1)(b + 1) = (ab + a + b + 1) > ab +100 (a + b) + 1 > 100 If the sum of the 2 sides > 99 then the area will > 100.

OpenStudy (anonymous):

Good job! Thanx a ton.......now how can I give you a medal. Since I am very new to this site I don't know how to give a medal yet.